A trip to Florida takes 29 1/4 hours. Assuming we are 1/4 of the way, how long has the trip taken so far?
Answer by askmemath(368)

Total time to travel to florida = 117/4
If we are 1/4 of the way = 117/4/4 = 117/16 = 7 5/16 has been travelled by us.